OpenStudy (anonymous):

Solve. 1/6r+1/4=1/3+3/4r

OpenStudy (anonymous):

\[\frac{1}{6}r+\frac{1}{4}=\frac{1}{3}+\frac{3}{4}r\] like that?

OpenStudy (anonymous):

multiply both sides by the least common multiple of 6 3 and 4, which is 12 that way, all the denominators will cancel and you will have whole numbers to work with

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[12\left(\frac{1}{6}r+\frac{1}{4}\right)=12\left(\frac{1}{3}+\frac{3}{4}r\right)\]

OpenStudy (anonymous):

6 goes in to 12 twice, 4 goes in to 12 3 times and 3 goes in to 12 4 times, so the denominators all go and you get \[2r+3=4+3\times 3r\] or \[2r+3=4+9r\] is that step ok or confusing?

OpenStudy (anonymous):

i'm trying to get it lol processing

OpenStudy (anonymous):

why did you multiply 12?

OpenStudy (anonymous):

ok i can write out in more detail if you like, but sometimes more detail is more confusing the idea is that you really don't want to do all the work with fractions, so if you multiply both sides by the least common multiple of the denominators they will go away

OpenStudy (anonymous):

in other words, find a number that 6, 3 and 4 go in to evenly

OpenStudy (anonymous):

it is pretty clear that they all divide 12 evenly, so if you multiply both sides of the equation by 12 there will be no more fractions

OpenStudy (anonymous):

oh! ok i got it

OpenStudy (anonymous):

you have to multiply everything by 12, which really means cancel as you go

OpenStudy (anonymous):

\[12\left(\frac{1}{6}r+\frac{1}{4}\right)=12\left(\frac{1}{3}+\frac{3}{4}r\right)\] \[12\times \frac{1}{6}r+12\times \frac{1}{4}=12\times \frac{1}{3}+12\times \frac{3}{4}r\] \[2r+3=4+9r\] you do the middle step in your head

OpenStudy (anonymous):

then it should be easier \[3=4+7r\] \[-1=7r\] \[-\frac{1}{7}=r\]

OpenStudy (anonymous):

hope all steps are clear you could work with the fractions but that would require adding and subtracting fractions, and dividing by fractions, which is more work

OpenStudy (anonymous):

It was a bit confusing before but you made it a lot easier thank you!

OpenStudy (anonymous):

yw